# Vision-Language-Action Models: Concepts, Progress, Applications and Challenges

### v1

@@ 3.1 Architectural Innovations in VLA Models
HybridVLA [110] 1. Vision Encoder: CLIP + DINOv2 2. Language Encoder: LLaMA-2 3. Action Decoder: Hybrid diffusion + autoregressive ensemble RT-X + synthetic task fusion Achieves robust control in complex multi-arm settings via dynamic ensemble; strong sim2real generalization
MoLe-VLA [213] 1. Vision Encoder: Multi-stage ViT with STAR router 2. Language Encoder: CogKD-enhanced transformer 3. Action Decoder: Sparse transformer with dynamic routing RLBench + real-world manipulation tasks Brain-inspired efficiency with 5.6x speedup; selective layer activation with high task success (+8%)
@@
UAV-VLA [150] 1. Vision Encoder: ViT for aerial imagery 2. Language Encoder: GPT for instruction parsing 3. Action Decoder: Transformer-based path planner Satellite + UAV imagery instructions Zero-shot aerial task planning; intuitive language grounding; scalable to large unmapped environments
@@
DexGraspVLA [219] 1. Vision Encoder: Object-centric spatial ViT 2. Language Encoder: Transformer with grasp sequence reasoning 3. Action Decoder: Diffusion controller for grasp pose generation Dexterous grasping benchmark (sim + real) 90%+ zero-shot success on diverse objects; excels at lighting, background variation, and unseen conditions
@@
GraspVLA [38] 1. Vision Encoder: Multi-view DINO-v2 and SigLIP fusion; InternLM2 language model 2. Language Encoder: VLM predicts bounding boxes and grasp poses 3. Action Decoder: Flow-matching based action expert via Progressive Action Generation (PAG) SynGrasp-1B (1B synthetic frames), GRIT (Internet grounding dataset) First synthetic-data-pretrained grasping VLA; enables sim-to-real generalization, robust grasp policy via PAG; supports zero-shot and few-shot generalization to long-tail object classes and human-centric preferences
Interleave-VLA [51] 1. Vision Encoder: InternVL2.5 and OWLv2 for open-vocabulary and image-text token integration 2. Language Encoder: Qwen2.5 for instruction parsing and visual-language verification 3. Action Decoder: Continuous action predictor adapted from OpenVLA and $\pi^{0}$ with diffusion-policy controller Open Interleaved X-Embodiment (210k episodes from 11 real-world datasets) First end-to-end VLA model for interleaved image-text instructions; improves out-of-domain generalization 2–3× and enables zero-shot execution from hand-drawn sketches and novel multimodal prompts
3.2 Training and Efficiency Advancements in Vision–Language–Action Models
@@
      Co-fine-tuning on massive vision–language corpora (e.g. LAION-5B) and robotic trajectory collections (e.g. Open X-Embodiment) aligns semantic understanding with motor skills. OpenVLA (7 B params) achieves a 16.5 % higher success rate than a 55 B-parameter RT-2 variant, demonstrating that co-fine-tuning yields strong generalization with fewer parameters [152, 175, 94].
@@
      Synthetic Data Generation via UniSim produces photorealistic scenes—including occlusions and dynamic lighting—to augment rare edge-case scenarios, improving model robustness in cluttered environments by over 20 % [200, 167].
@@
      Compressed Action Tokens (FAST) and Parallel Decoding in dual-system frameworks (e.g. Groot N1) yield 2.5× faster policy steps, achieving sub-5 ms latencies at a modest cost to trajectory smoothness [13, 161].
    • (b)
      Hardware-Aware Optimizations—including tensor-core quantization and pipelined attention kernels—shrink runtime memory footprints below 8 GB and enable real-time inference on embedded GPUs [93].
@@
3.3 Parameter‐Efficient Methods and Acceleration Techniques in VLA Models
Building on advances in data‐efficient training, recent work has focused on reducing the parameter footprint and improving inference speed of VLA models—critical for deployment on resource‐constrained robotic platforms.
@@ 3.3 Parameter‐Efficient Methods and Acceleration Techniques in VLA Models
    Low‐Rank Adaptation (LoRA). LoRA injects small trainable rank‐decomposition matrices into frozen transformer layers, enabling fine‐tuning of billion‐parameter VLAs with only a few million additional weights. In OpenVLA, LoRA adapters (20 M parameters) tuned a 7 B‐parameter backbone on commodity GPUs in under 24 h, cutting GPU compute by 70 % compared to full backpropagation [72, 94]. Crucially, LoRA‐adapted models retain their high‐level language grounding and visual reasoning capabilities while adapting to new robotic manipulation tasks (e.g. novel object shapes), making large VLAs accessible to labs without supercomputing resources.
@@
    Quantization. Reducing weight precision to 8‐bit integers (INT8) shrinks model size by half and doubles on‐chip throughput. OpenVLA experiments show that INT8 quantization on Jetson Orin maintains 97 % of full‐precision task success across pick‐and‐place benchmarks, with only a 5 % drop in fine‐grained dexterity tasks [152, 94]. Complementary methods such as post‐training quantization with per‐channel calibration further minimize accuracy loss in high‐dynamic‐range sensor inputs [128]. These optimizations allow continuous control loops at 30 Hz on 50 W edge modules.
@@
    Model Pruning. Structured pruning removes entire attention heads or feed‐forward sublayers identified as redundant. While less explored in VLA than in pure vision or language models, early studies on Diffusion Policy demonstrate that pruning up to 20 % of ConvNet‐based vision encoders yields negligible performance degradation in grasp stability [34]. Similar schemes applied to transformer‐based VLAs (e.g. RDT‐1B) can reduce memory footprint by 25 % with under 2 % drop in task success, paving the way for sub‐4 GB deployments [112, 102].
@@
    Compressed Action Tokenization (FAST). FAST reformulates continuous action outputs as frequency‐domain tokens, compressing long control sequences into concise descriptors. The Pi‐0 Fast variant achieved 15× faster inference with a 300 M‐parameter diffusion head by tokenizing 1000 ms action windows into 16 discrete tokens, enabling 200 Hz policy rates on desktop GPUs [133]. This approach trades minimal trajectory granularity for large speedups, suited for high‐frequency control in dynamic tasks like bimanual assembly.
@@
    Parallel Decoding and Action Chunking. Autoregressive VLAs traditionally decode actions token by token, incurring sequential latency. Parallel decoding architectures (e.g. in Groot N1) decode groups of spatial–temporal tokens concurrently, achieving a 2.5× reduction in end‐to‐end latency on 7‐DoF arms at 100 Hz, with less than 3 mm positional error increase [13, 161]. Action chunking further abstracts multi‐step routines into single tokens (e.g. “pick‐and‐place‐cup”), cutting inference steps by up to 40 % in long‐horizon tasks like kitchen workflows [86].
@@
    Reinforcement Learning–Supervised Hybrid Training. The iRe‐VLA framework alternates between reinforcement learning (RL) in simulation and supervised fine‐tuning on human demonstrations to stabilize policy updates. By leveraging Direct Preference Optimization (DPO) to shape reward models and Conservative Q‐Learning to avoid extrapolation error, iRe‐VLA reduces sample complexity by 60 % versus pure RL, while maintaining the semantic fidelity imparted by language‐conditioned priors [123, 65]. This hybrid approach yields robust policies for tasks with sparse feedback, such as dynamic obstacle avoidance.
  7. 7.
    Hardware‐Aware Optimizations. Compiler‐level graph rewrites and kernel fusion (e.g. via NVIDIA TensorRT‐LLM) exploit target hardware features—tensor cores, fused attention, and pipelined memory transfers—to accelerate both transformer inference and diffusion sampling. In OpenVLA‐OFT, such optimizations reduced inference latency by 30 % on RTX A2000 GPUs and lowered energy per inference by 25 % compared to standard PyTorch execution [93]. This makes real‐time VLAs feasible on mobile robots and drones with strict power budgets.
Discussion. Parameter‐efficient adaptation and inference acceleration techniques collectively democratize VLA deployment:
@@
    LoRA and quantization empower smaller labs to fine‐tune and operate billion‐parameter VLAs on consumer‐grade hardware, unlocking cutting‐edge semantic understanding for robots [72, 94].
@@ 3.4 Applications of Vision-Language-Action Models
VLA models are rapidly emerging as foundational building blocks for embodied intelligence, integrating perception, natural language understanding, and motor control within a unified architecture. By encoding visual and linguistic modalities into shared semantic spaces and generating contextually grounded actions, VLA models enable seamless interaction between agents and their environments [102, 220]. This multimodal capacity has positioned VLAs as transformative agents across a wide spectrum of real-world applications. In humanoid robotics, systems like Helix and RoboNurse-VLA combine vision, language, and dexterous manipulation to assist with domestic tasks and surgical operations, demonstrating real-time reasoning and safety-aware control [103, 186]. In autonomous vehicles, models such as OpenDriveVLA and ORION process dynamic visual streams and natural language instructions to make transparent, adaptive driving decisions in complex urban environments [56, 220]. Industrial deployments leverage VLA architectures for high-precision assembly, inspection, and collaborative manufacturing [102]. In agriculture, VLA-powered robotic systems enable vision-guided fruit harvesting, plant monitoring, and anomaly detection, reducing labor dependency and increasing sustainability. Furthermore, recent advances in interactive augmented reality systems utilize VLA models for real-time, language-conditioned spatial navigation, guiding users in indoor and outdoor settings based on voice or visual cues [150, 59]. Across these domains, VLAs offer a unified framework for robust, adaptable, and semantically aligned task execution, marking a pivotal shift toward embodied generalist agents.
@@
Table 3 in the appendix shows the recent VLA models by summarizing their methodologies, application domains, and key innovations.
@@
Figure 11: Mind‐map of application domains for Vision–Language–Action models.
Table 3: Comparison of VLA methodologies, application areas, and innovations. This comprehensive table compares cutting-edge VLA models by summarizing their methodologies, application domains, and key innovations.
Reference & Year VLA Methodology VLA Application Area Strength and Key Innovation
CogACT [102] & 2024 Componentized VLA with specialized action module using diffusion transformers Industrial robotics, language-guided manipulation Robust action modeling, rapid adaptation, strong generalization, much higher task success rates
VLATest [182] & 2024 Automated framework for large-scale VLA model testing in manipulation Robotic manipulation: benchmarking VLA robustness and reliability Diverse scene generation, multi-model/task evaluation, reveals robustness gaps, guides VLA improvement
NaVILA [32] & 2024 Two-level VLA: high-level vision-language generates mid-level nav commands, RL locomotion executes Legged robot navigation via natural language in cluttered, real-world scenes Modular mid/low-level split, strong generalization, 88% real-world success, robust to diverse terrains
RoboNurse-VLA [103] & 2024 SAM 2 vision, Llama 2 language, real-time voice-to-action pipeline Surgical assistance: precise instrument grasp and handover in OR Accurate, real-time handover, robust to unseen tools, excels at complex, dynamic surgical tasks
Mobility VLA [35] & 2024 Hierarchical VLA: long-context VLM for multimodal goal localization, topological graph for low-level navigation Multimodal instruction navigation with demonstration tours (MINT) in real-world environments High success on complex language+image tasks, robust to novel queries, leverages demonstration videos, scalable to large spaces
CoVLA [5] & 2025 CLIP for vision, Llama-2 for language, trajectory prediction for action Autonomous driving, dataset for VLA model training Large-scale, richly annotated dataset; enables interpretable scene understanding and robust path planning
OpenDriveVLA [220] & 2025 Hierarchical alignment of 2D/3D visual tokens and language embeddings; autoregressive agent-env-ego modeling End-to-end autonomous driving Unified semantic space, dynamic interaction modeling, state-of-the-art planning and QA performance
ORION [56] & 2025 QT-Former for history context, LLM for reasoning, generative planner for trajectory prediction Holistic E2E autonomous driving Aligns reasoning and action spaces, unified optimization for VQA and planning, superior closed-loop results
QUAR-VLA [43] & 2025 QUART model fuses vision and language for action generation Quadruped robots: navigation, manipulation, whole-body tasks Tight vision-language-action integration, fine-grained instruction alignment, strong sim-to-real generalization
TinyVLA [186] & 2025 Compact VLA with fast multimodal backbone, diffusion policy decoder Robotic manipulation: fast, data-efficient visuomotor control No pre-training needed, faster inference, strong generalization, outperforms OpenVLA on efficiency and accuracy
UAV-VLA [150] & 2025 Modular VLA: GPT for goal extraction, VLM for object search, GPT for action generation Large-scale UAV mission planning from natural language and satellite imagery Efficient flight path/action plan generation, no prior training, intuitive human-UAV interaction, benchmarked performance
Bi-VLA [59] & 2025 Multimodal transformer links vision, language, and bimanual action modules Bimanual dexterous manipulation for household tasks Accurate code/action generation, high adaptability, strong vision-language-action integration, robust real-world performance
ChatVLA [221] & 2025 Phased Alignment Training, Mixture-of-Experts for vision-language-action integration Unified multimodal understanding and robot control Minimizes forgetting/interference, excels at VQA and manipulation, efficient, outperforms SOTA VLA models
RoboMamba [111] & 2025 Mamba-based VLA: vision encoder co-trained with SSM for reasoning and SE(3) action Robotic reasoning and manipulation, efficient pose prediction Linear-complexity inference, minimal fine-tuning, fast and accurate reasoning and manipulation, SOTA efficiency
OTTER [75] & 2025 Text-aware visual feature extraction with frozen pre-trained VLMs Robotic manipulation: zero-shot generalization to novel tasks Preserves semantic alignment, no VLM fine-tuning, task-relevant feature selection, strong zero-shot performance
PointVLA [96] & 2025 Injects 3D point cloud features into frozen pre-trained VLA via modular skip-blocks Robotic manipulation: spatial reasoning, few-shot, long-horizon tasks No retraining, preserves 2D knowledge, strong 3D spatial reasoning, excels at few-shot and dynamic tasks
VLA-Cache [195] & 2025 Adaptive token caching with selective reuse of static visual tokens Robotic manipulation: efficient, real-time VLA inference 40–50% faster, minimal accuracy loss, dynamic layer-wise token reuse, practical for real-world robots
CombatVLA [29] & 2025 Trains on video-action AoT sequences, integrates truncated AoT for fast inference 3D ARPGs: real-time combat understanding and tactical action 50x faster inference, outperforms all baselines, surpasses human success rate, strong tactical reasoning
HybridVLA [110] & 2025 Unified LLM with collaborative diffusion and autoregressive action policies Robotic manipulation: single-arm, dual-arm, diverse real/sim tasks Adaptive action ensemble, robust control, strong generalization, outperforms SOTA on complex manipulations
NORA [79] & 2025 3B-parameter VLA using Qwen-2.5-VL-3B backbone and FAST+ tokenizer Generalist embodied robotics: efficient real-world and simulated task execution Low computational overhead, strong visual reasoning, fast action decoding, outperforms larger VLA models
SpatialVLA [136] & 2025 Ego3D Position Encoding and Adaptive Action Grids for spatially-aware VLA Generalist robot manipulation: cross-robot, multi-task, zero-shot control 3D spatial integration, adaptive action discretization, strong generalization and transfer, open-sourced code
MoLe-VLA [213] & 2025 Mixture-of-Layers with dynamic layer-skipping via STAR router and CogKD Efficient robot manipulation: RLBench and real-world tasks Selective layer activation, 5.6x faster, preserves cognition, +8% mean success, brain-inspired efficiency
JARVIS-VLA [101] & 2025 Post-trains large VLMs with visual-language guidance and action head for keyboard/mouse control Visual games (Minecraft): 1k+ tasks, open-world, instruction following Self-supervised post-training, 40%+ over baselines, strong world knowledge, state-of-the-art generalization, open-sourced
UP-VLA [209] & 2025 Unified VLA with joint multi-modal understanding and future prediction objectives Embodied agents: manipulation tasks, precise spatial reasoning Captures both high-level semantics and low-level spatial dynamics, 33% better on Calvin ABC-D, excels at real-world tasks needing fine spatial control
Shake-VLA [92] & 2025 Modular VLA system with vision, speech-to-text, RAG, anomaly detection, and bimanual arms Bimanual robotic cocktail mixing: ingredient detection, recipe adaptation, liquid measurement 100% task success, robust in noisy/cluttered settings, accurate ingredient handling, flexible recipe adaptation, real-world deployment
MoRE [214] & 2025 Sparse Mixture-of-Experts (MoE) with LoRA modules, RL-based Q-function training Quadruped robots: multi-task navigation, locomotion, and manipulation Scalable RL fine-tuning on mixed-quality data, strong multi-task and OOD generalization, outperforms baselines in sim and real-world
DexGraspVLA [219] & 2025 Hierarchical VLA: pre-trained vision-language planner + diffusion-based low-level controller General dexterous grasping: robust across diverse objects, lighting, and backgrounds Iterative domain-invariant representation, strong zero-shot generalization, 90%+ success on unseen scenarios, consistent performance across variations
DexVLA [185] & 2025 Plug-in billion-param diffusion action expert, embodiment curriculum learning General robot control: single-arm, bimanual, dexterous hand, long-horizon tasks Cross-embodiment action modeling, efficient curriculum training, rapid adaptation, SOTA on complex tasks without task-specific tuning

### v2

@@ 3.1 Architectural Innovations in VLA Models
HybridVLA [142] CLIP + DINOv2 / LLaMA-2 / hybrid diffusion + autoregressive ensemble RT-X + synthetic fusion Dynamic ensemble improves robustness in multi-arm settings and supports stronger sim-to-real generalization.
[…]
@@
UAV-VLA [191] ViT for aerial imagery / GPT instruction parsing / transformer path planner Satellite + UAV imagery instructions Zero-shot aerial task planning with scalable language grounding for large unmapped environments.
@@
DexGraspVLA [291] Object-centric spatial ViT / transformer grasp reasoning / diffusion grasp controller Dexterous grasping benchmark (sim+real) Reported >>90% zero-shot success across diverse objects; robust to lighting/background variation and unseen conditions.
@@
GraspVLA [46] Multi-view DINOv2 + SigLIP / VLM predicts boxes+grasps / flow-matching action expert (PAG) SynGrasp-1B; GRIT Synthetic-pretrained grasping VLA; improves sim-to-real transfer and supports zero-/few-shot generalization to long-tail object classes and preferences.
[…]
Long-VLA [64] End-to-end VLA for long-horizon tasks / phase-aware input masking + transformer policy (subtask phase segmentation) Long-horizon multi-step robot manipulation demonstrations (task sequences) Targets long-horizon execution by explicitly separating “moving” vs. “interaction” phases, improving subtask compatibility and robustness over extended task horizons.
RetoVLA [123] VLM-based policy / reuses register tokens as spatial context for action prediction Real-robot manipulation on a 7-DoF arm + task-specific demonstrations Lightweight spatial-reasoning upgrade by repurposing register tokens; reported sizable success-rate gains on complex manipulation with minimal architectural overhead.
Vlaser [256] VLM-to-VLA pipeline / synergistic embodied reasoning + policy learning (reasoning-aware VLA fine-tuning) Vlaser-6M embodied reasoning dataset + VLA fine-tuning data (robot demonstrations) Bridges embodied reasoning and control: strong performance across embodied grounding/QA/planning while improving transfer to policy learning under domain shift.
Discrete Diffusion VLA [138] Single-transformer VLA / discretized action chunks + discrete diffusion refinement (CE training; remasking) LIBERO + SimplerEnv (Fractal/Bridge) style VLA benchmarks Unifies diffusion-style refinement with discrete token interfaces: adaptive decoding order, error correction via remasking, and reduced autoregressive bottlenecks with strong benchmark success rates.
Being-H0 [152] Dexterous VLA pretrained from human videos / explicit hand-motion modeling + VL grounding for action Large-scale human manipulation videos (egocentric/third-person) + transfer to robot control Scales dexterous manipulation by leveraging diverse human video data; improves generalization to novel tasks/scenes compared with small teleop-only robot datasets.
EgoVLA [257] VLM pretraining on egocentric human manipulation / unified human–robot action space + robot fine-tuning Large-scale egocentric human videos + small set of robot demonstrations Uses abundant egocentric videos for scalable pretraining, then aligns embodiments via unified action space to enable practical robot transfer with limited robot data.
StereoVLA [47] Stereo-enhanced VLA / geometric-semantic fusion from stereo pairs (+ auxiliary depth cues) for action decoding Stereo robot manipulation demonstrations (stereo RGB with language-conditioned actions) Explicitly exploits stereo geometry to improve spatial precision (depth-sensitive grasping/manipulation) and robustness to viewpoint/camera variations.
EfficientVLA [260] Training-free VLA acceleration/compression / structured redundancy removal across VLA pipeline (token/interface optimizations) Applies to existing VLAs (evaluation on standard manipulation benchmarks; e.g., SIMPLER-style settings) Practical deployability: speeds up and reduces compute/memory of large VLA policies with minimal accuracy drop, enabling closer-to-real-time inference on constrained hardware.
3.2 Training Efficiency Advancements in Vision–Language–Action Models
@@
      Co-fine-tuning on massive vision–language corpora (e.g. LAION-5B) and robotic trajectory collections (e.g. Open X-Embodiment) aligns semantic understanding with motor skills. OpenVLA (7 B parameters) achieves a 16.5 % higher success rate than a 55 B-parameter RT-2 variant, demonstrating that co-fine-tuning yields strong generalization with fewer parameters [194, 227, 122] compared to compared to scaling model size alone.
@@
      Synthetic Data Generation via UniSim produces photorealistic scenes including occlusions and dynamic lighting to augment rare edge-case scenarios, improving model robustness in cluttered environments by over 20 % [264, 218].
@@
[…]
@@
3.3 Parameter-Efficient Methods and Acceleration Techniques in VLA Models
Beyond data-efficient learning strategies, a complementary line of research in VLA models focuses on reducing model size, memory footprint, and inference latency to enable deployment on real robotic platforms with limited computational and power resources. Unlike training-centric efficiency methods discussed earlier, the techniques in this subsection primarily target parameter efficiency at adaptation time and runtime acceleration during policy inference.
@@ 3.3 Parameter-Efficient Methods and Acceleration Techniques in VLA Models
    Parameter-Efficient Adaptation via Low-Rank Modules: Rather than full fine-tuning, many VLAs adopt parameter-efficient adaptation mechanisms such as Low-Rank Adaptation (LoRA), which was introduced earlier in the context of training efficiency. In this section, we emphasize its role in reducing the effective parameter footprint during deployment. For example, OpenVLA employs lightweight LoRA adapters (approximately 20 M parameters) on top of a frozen 7 B-parameter backbone, allowing task adaptation with minimal memory overhead and avoiding duplication of full model weights across tasks [93, 122]. This design enables multiple task-specialized policies to coexist on resource-constrained systems while preserving the general-purpose visual–semantic representations learned during pretraining.
@@
    Quantization for Edge Deployment: Model quantization reduces numerical precision to improve inference throughput and memory efficiency. Experiments with OpenVLA show that INT8 quantization on embedded platforms such as NVIDIA Jetson Orin preserves approximately 97 % of full-precision task success on pick-and-place benchmarks, with only minor degradation in fine-grained dexterous manipulation [194, 122]. Post-training quantization and per-channel calibration further mitigate accuracy loss under high-dynamic-range sensory inputs [164]. These optimizations enable sustained control frequencies of up to 30 Hz within strict power budgets typical of mobile robots.
@@
    Model Pruning and Architectural Slimming: Structured pruning removes redundant architectural components such as attention heads or feed-forward sublayers to reduce memory and compute requirements. Although less explored in VLAs than in standalone vision or language models, early studies on diffusion-based visuomotor policies indicate that pruning up to 20 % of convolutional vision encoders results in negligible degradation in grasp stability [40]. Similar pruning strategies applied to transformer-based VLAs (e.g. RDT-1B) can reduce memory footprint by approximately 25 % with less than 2 % drop in task success, enabling sub-4 GB deployments [145, 131].
@@
[…]
@@
[…]
@@
    Hardware-Aware Compilation and Runtime Optimization: Finally, hardware-aware optimizations leverage compiler-level graph rewrites, kernel fusion, and accelerator-specific primitives to maximize throughput. Frameworks such as TensorRT-LLM exploit tensor cores, fused attention kernels, and pipelined memory transfers to accelerate both transformer inference and diffusion sampling. In OpenVLA-OFT, such optimizations reduce inference latency by approximately 30 % and lower energy consumption per inference by 25 % on RTX-class GPUs compared to standard PyTorch execution [121]. These system-level optimizations are critical for deploying real-time VLAs on mobile robots, aerial platforms, and humanoid systems with strict power constraints.
Discussion: Parameter‐efficient adaptation and inference acceleration techniques collectively democratize VLA deployment:
@@
    LoRA and quantization empower smaller labs and research/development programs to fine‐tune and operate billion‐parameter VLAs on consumer‐grade hardware, unlocking cutting‐edge semantic understanding for robots [93, 122].
@@ 3.4 Applications of Vision-Language-Action Models
VLA models are rapidly emerging as foundational building blocks for embodied intelligence, integrating perception, natural language understanding, and motor control within a unified architecture. By encoding visual and linguistic modalities into shared semantic spaces and generating contextually grounded actions, VLA models enable seamless interaction between agents and their environments [131, 293]. By encoding visual and linguistic modalities into shared semantic spaces and generating contextually grounded actions, VLA models enable seamless interaction between agents and their environment. This multimodal capacity has positioned VLAs as transformative agents across a wide spectrum of real-world applications. In humanoid robotics, systems like Helix and RoboNurse-VLA combine vision, language, and dexterous manipulation to assist with domestic tasks and surgical operations, demonstrating real-time reasoning and safety-aware control [132, 242]. In autonomous vehicles, models such as OpenDriveVLA and ORION process dynamic visual streams and natural language instructions to make transparent, adaptive driving decisions in complex urban environments [69, 293]. Industrial deployments leverage VLA architectures for high-precision assembly, inspection, and collaborative manufacturing [131]. In agriculture, VLA-powered robotic systems could enable vision-guided fruit harvesting, plant monitoring, and anomaly detection, reducing labor dependency and increasing sustainability. Furthermore, recent advances in interactive augmented reality systems utilize VLA models for real-time, language-conditioned spatial navigation, guiding users in indoor and outdoor settings based on voice or visual cues [191, 74]. Across these domains, VLAs offer a unified framework for robust, adaptable, and semantically aligned task execution, marking a pivotal shift toward embodied generalist agents.
@@
Table 3 shows the recent VLA models by summarizing their methodologies, application domains, and key innovations.
Table 3: Comparison of representative Vision-Language-Action (VLA) methodologies, application areas, and key innovations.
Reference (Year) VLA methodology Application area Strength / key innovation
CogACT [131] (2024) Modular VLA with a specialized action module based on diffusion transformers. Industrial robotics; language-guided manipulation. Robust action modeling with fast adaptation and strong generalization, improving task success in diverse industrial settings.
VLATest [237] (2024) Automated large-scale testing framework for VLA manipulation evaluation. Robotic manipulation benchmarking (robustness/reliability). Systematic multi-scene and multi-task evaluation that exposes robustness gaps and supports targeted VLA improvements.
NaVILA [38] (2024) Two-level VLA: high-level vision-language generates mid-level navigation commands; RL locomotion executes. Legged navigation from natural language in cluttered real-world scenes. Modular high/low-level split with strong generalization and high real-world success across challenging terrains.
RoboNurse-VLA [132] (2024) Vision module (SAM2) + language module (Llama2) with real-time voice-to-action pipeline. Surgical assistance (instrument grasp and handover). Accurate real-time assistance, robustness to unseen tools, and stable performance in dynamic operating room conditions.
Mobility VLA [42] (2024) Hierarchical VLA with long-context VLM for goal localization and topological graph navigation. Multimodal instruction navigation with demonstration tours (MINT). Leverages demonstrations for scalable navigation in large spaces; robust to novel language+image queries.
CoVLA [5] (2025) CLIP-based vision, Llama-2 language, trajectory prediction for action. Autonomous driving (dataset + VLA training). Large-scale richly annotated dataset enabling interpretable scene understanding and robust path planning.
OpenDriveVLA [293] (2025) Hierarchical alignment of 2D/3D visual tokens and language embeddings; autoregressive agent–environment–ego modeling. End-to-end autonomous driving. Unified semantic space and interaction modeling that improves planning and QA performance in complex traffic scenes.
ORION [69] (2025) QT-Former for history context, LLM reasoning, generative planner for trajectory prediction. Holistic end-to-end autonomous driving. Aligns reasoning and action spaces with unified optimization for VQA and planning, yielding stronger closed-loop performance.
QUAR-VLA [54] (2025) QUART-based fusion of vision and language for action generation. Quadruped navigation, manipulation, whole-body tasks. Tight VL-A coupling and instruction alignment with strong sim-to-real generalization for legged platforms.
TinyVLA [242] (2025) Compact multimodal backbone with diffusion-policy decoder. Fast, data-efficient manipulation control. Reduced inference cost with strong generalization; improves efficiency without requiring heavy pre-training.
UAV-VLA [191] (2025) Modular pipeline: GPT for goal extraction, VLM for object search, GPT for action generation. UAV mission planning from language + satellite imagery. Efficient flight/action planning with minimal prior training; supports intuitive human–UAV interaction and benchmarking.
Bi-VLA [74] (2025) Multimodal transformer linking vision, language, and bimanual action modules. Bimanual household manipulation. High adaptability and robust real-world bimanual execution through explicit action-module integration.
ChatVLA [295] (2025) Phased alignment training with Mixture-of-Experts for VL-A integration. Unified multimodal understanding and robot control. Reduces forgetting/interference, improving both VQA and manipulation with efficient specialization.
RoboMamba [143] (2025) Mamba/SSM-based VLA with co-trained vision encoder and SE(3) action modeling. Efficient robotic reasoning and manipulation. Linear-complexity inference and minimal fine-tuning enable fast pose-aware reasoning and efficient control.
OTTER [97] (2025) Text-aware feature extraction using frozen pre-trained VLMs. Manipulation with zero-shot generalization. Preserves VL semantic alignment without VLM fine-tuning via task-relevant feature selection for strong zero-shot transfer.
PointVLA [125] (2025) Injects 3D point-cloud features into frozen VLA via modular skip-blocks. Spatial reasoning; few-shot and long-horizon manipulation. Adds 3D geometry while preserving 2D knowledge, improving spatial grounding without full retraining.
VLA-Cache [252] (2025) Adaptive token caching with selective reuse of static visual tokens. Real-time efficient manipulation inference. Layer-wise token reuse yields major speedups with minimal accuracy loss, practical for on-robot deployment.
CombatVLA [34] (2025) Video-action AoT training with truncated AoT for fast inference. Real-time combat decision-making in 3D games. Large inference acceleration with improved tactical reasoning and high success in real-time interactive environments.
HybridVLA [142] (2025) Unified LLM with collaborative diffusion and autoregressive action policies. Single-/dual-arm manipulation across sim and real tasks. Adaptive action ensembling improves robustness and generalization on complex manipulations.
NORA [103] (2025) 3B-parameter VLA using Qwen-2.5-VL-3B backbone with FAST+ tokenizer. Generalist embodied robotics (sim + real). Low overhead with strong visual reasoning and fast action decoding; competitive with larger VLAs.
SpatialVLA [175] (2025) Ego3D position encoding and adaptive action grids for spatially-aware VLA. Cross-robot, multi-task, zero-shot manipulation. Explicit 3D spatial integration and adaptive discretization improve transfer and generalization; open-sourced.
MoLe-VLA [283] (2025) Mixture-of-Layers with dynamic layer-skipping via router and distillation. Efficient manipulation on RLBench + real robots. Selective layer activation provides strong speedups while maintaining cognition and improving success.
JARVIS-VLA [130] (2025) Post-trained large VLMs with VL guidance and an action head for keyboard/mouse control. Open-world visual games (e.g., Minecraft), 1k+ tasks. Self-supervised post-training yields strong generalization and world knowledge; open-sourced task suite and models.
UP-VLA [279] (2025) Unified VLA with joint multimodal understanding and future prediction objectives. Embodied manipulation with precise spatial reasoning. Joint semantics+dynamics learning improves performance on long-horizon tasks requiring fine spatial control.
Shake-VLA [120] (2025) Modular stack with vision, speech-to-text, RAG, anomaly detection, and bimanual arms. Bimanual cocktail mixing in clutter/noise. Robust end-to-end deployment with reliable ingredient handling, recipe adaptation, and real-world task completion.
MoRE [285] (2025) Sparse MoE with LoRA modules and RL-based Q-function training. Quadruped multi-task locomotion, navigation, manipulation. Scalable fine-tuning on mixed-quality data with strong multi-task and OOD generalization in sim and real.
DexGraspVLA [291] (2025) Hierarchical planner (pre-trained VL) + diffusion low-level controller. General dexterous grasping in diverse conditions. Domain-invariant representations support strong zero-shot generalization across objects, lighting, and backgrounds.
DexVLA [241] (2025) Plug-in diffusion action expert with embodiment curriculum learning. General robot control: single-arm, bimanual, dexterous hand. Cross-embodiment action modeling and fast adaptation enable strong long-horizon performance without task-specific tuning.
@@
Figure 11: Mind-map of application domains for Vision–Language–Action models, with Humanoid Robotics positioned at the top and remaining domains arranged clockwise to match the order of discussion in this section.